# MACI: Multi-Agent Collaborative Intelligence for Robust Reasoning and Temporal Planning

### v1

@@ 4.2.2 State Space and Agent Design
Common agents are designed to handle core planning and validation tasks that are universally applicable across domains. These agents provide foundational capabilities that ensure the consistency, feasibility, and robustness of planning systems, regardless of the specific application. Their roles align with the general state space dimensions of Who, Where, When, What, and Why, ensuring that all essential aspects of planning are addressed. For a list of common agents, please refer to Appendix D.2.
These agents work collaboratively to ensure robust planning and validation across diverse domains. Although the list of common agents is not exhaustive, it provides a strong foundation to address most planning tasks. New common agents can be introduced as the need arises, provided they offer capabilities that can be reused across multiple domains. For example, an Adaptation Agent could be added to dynamically adjust constraints in response to evolving task environments.
Specialized Agents
Specialized agents extend the capabilities of MACI by addressing domain-specific requirements that are not generalizable across all tasks. These agents are designed to implement detailed knowledge and expertise tailored to specific applications, ensuring that MACI remains adaptable and effective for diverse planning challenges. Specialized agents typically focus on tasks requiring unique constraint handling, optimization, or analysis.
Examples of specialized agents include the following.
  • Travel Planning Agent: Optimizes itineraries, manages transportation options, and ensures seamless transitions between locations.
  • Financial Risk Agent: Monitors market trends, assesses risks, and validates investment strategies.
  • Healthcare Resource Agent: Allocates medical supplies, schedules staff, and optimizes patient care pathways.
  • Supply Chain Agent: Manages logistics, tracks inventory, and optimizes distribution routes.
Specialized agents adhere to architectural principles that maintain consistency and integration within MACI:
  • Restricted Context Windows: Each agent operates within a limited context to ensure efficiency and focus ($w\leq 1k$ tokens).
  • Protocol Buffer Interfaces: Standardized interfaces enable seamless communication between agents.
  • Coordination with Common Agents: Specialized agents leverage the capabilities of common agents for validation, compliance, and constraint management.
  • Domain-Specific Constraints: Specialized agents are equipped to handle unique requirements and challenges specific to their domain.
4.2.3 Agent Coordination and State & Constraint Management
MACI ensures seamless coordination between agents across state dimensions. For example, in travel planning, when a Schedule Agent detects a flight connection constraint, it triggers updates to Location, Resource, and Role Manager Agents. These updates propagate through a state transition protocol, where the proposed state change is validated against global and local constraints:
$T:S_{t}\rightarrow S_{t+1}\text{ subject to }C_{\text{global}}\land C_{\text{local}}$
4.2.4 Validation Protocols
The validation agent ensures state consistency by decomposing global constraints into dimension-specific checks. For example, when booking a flight, the temporal agent validates the temporal constraint, the spatial feasibility is checked by the location agent, and the availability of seats is validated by the resource agent. Common sense agents identify overlooked constraints, such as luggage claiming time, inter-terminal walking time, and rental car pickup time. If validation fails, the system triggers a recovery process to generate alternative solutions. Validation protocols can be updated when unexpected events occur to improve retrospective planning.
4.2.5 Other Considerations
Due to space limitation, we document in Appendix C additional considerations including cross-domain generalization, enhancement and integration of common sense knowledge, scalability, and a list of future enhancements when more applications will be supported.
@@ 4.3 Meta-Planner: Planning a Planner to Plan
The meta-planner $\mathcal{MP}$ operates as a higher-level planner that constructs other planners, which are actionable workflows for specific tasks.
@@ 4.3.1 The Meta-Planner Algorithm
$\mathcal{MP}$ is specified in Algorithm #1. A list of all symbols used in the specifications is provided in Table 20 in Appendix D.
A workflow network $\mathbf{W}=(\mathcal{N},\mathcal{E})$, represented as a directed graph, consists of roles as nodes and their dependencies as edges. The nodes $n\in\mathcal{N}$ denote the roles performed by human actors (e.g. driver, cook, supervisor), while the edges $e_{ij}\in\mathcal{E}$ between the nodes $n_{i}$ and $n_{j}$ capture the monitored dependencies (e.g., temporal, spatial constraints, data dependencies and safety requirements for supervision).
$\mathbf{W}=(\mathcal{N},\mathcal{E}),\text{ where }\mathcal{N}=A^{*}_{n}\text{and }\mathcal{E}=A^{*}_{e}.\vspace{-.1in}$ (1)
@@
  1. $\mathcal{N}$: Extract roles & capabilities from $\mathcal{O}$ (Eq. 2)
  2. $\mathcal{E}$: Identify node dependencies from $\mathcal{C_{E}}$ (Eq. 3)
@@
  3. Assign role agents: $\forall n\in\mathcal{N}$, select $\alpha_{n}\in\mathbf{A_{n}}$ (Eq. 4)
  4. Assign edge agents: $\forall e_{ij}\in\mathcal{E}$, $\alpha_{ij}\in\mathbf{A_{e}}$ (Eq. 5)
@@
        Update role-person mappings via $f_{\text{role}}(n,\mathcal{P})$
@@
  return $\mathbf{W^{*}}=\mathbf{W{current}}$
$\mathcal{MP}$ constructs a planner in three phases. First, it builds the network structure by identifying roles and their dependencies, forming the foundation of the workflow.
Second, it assigns appropriate agents to manage both roles and dependencies. Each node $n\in\mathcal{N}$ is managed by a designated role agent $\alpha_{n}$. Similarly, each edge $e_{ij}\in\mathcal{E}$, representing a dependency between nodes, is managed by an agent $\alpha_{ij}$, which connects the node agents $\alpha_{i}$ and $\alpha_{j}$. These edges are enriched with metadata that may include details such as data types, communication protocols, timing constraints, and security requirements.
Finally, the generated planner iteratively refines the workflow. This process involves optimizing role assignments and continuously monitoring dependencies until the validation score stabilizes, ensuring a well-validated workflow.
4.3.2 Meta-Planning Components
Let us clearly distinguish between roles (requiring human qualifications) and monitoring tasks (requiring capabilities).
@@ 4.3.2 Meta-Planning Components
$\mathcal{MP}$ analyzes objectives $\mathcal{O}$ to identify required roles and their qualifications:
@@
where $n_{i}$ represents a role and $q_{i}$ its required qualifications. For example, $(n_{\text{driver}},q_{\text{license}})$ or $(n_{\text{cook}},q_{\text{experience}})$.
@@
The framework maintains a global constraint set $C$ composed of:
@@
where $C_{E}$ represents explicit constraints from problem specification, $C_{I}$ denotes implicit constraints identified by Common Sense Agent, and $C_{D}$ represents derived constraints from agent interactions.
@@
Two types of agents, node and edge, are selected from the repository $\mathbf{A}$:
1. Node Monitoring Agents:
@@
These agents verify role qualifications and manage people-role assignments and transitions.
2. Edge Monitoring Agents:
@@
These agents monitor dependencies between roles, where $c_{j}$ represents constraint types (temporal, spatial, etc.).
$\mathcal{MP}$ iteratively refines the network by validating workflow constraints and performance and replacing or adding agents, i.e., checking and adding nodes and edges.
@@ 4.4 Workflow Execution Framework
The workflow $\mathbf{W^{*}}$ requires a runtime system for execution. In this work, we evaluate $\mathbf{W^{*}}$ by entering $\mathbf{W^{*}}$ along with the problem statement into an LLM (DeepSeek and GPT4o). A key limitation is that the feedback loop for refining $\mathbf{W^{*}}$ must be performed manually.
5 Evaluating $\mathcal{MP}$’s Planning Capabilities
To assess $\mathcal{MP}$’s performance and adaptability, we adopted a dual-approach experimental structure. The first experiment uses the Traveling Salesperson Problem (TSP) to validate $\mathcal{MP}$’s optimization capabilities. The second involves the Thanksgiving Dinner Planning problem, showcasing $\mathcal{MP}$’s ability to handle complex, real-world challenges with cross-thread dependencies and dynamic adaptability.
Traveling Salesperson Problem (TSP)
The TSP experiment benchmarks $\mathcal{MP}$ against planners (Claude, DeepSeek, GPT4o) and their $\mathcal{MP}$ integrations. Metrics include: 1) solving quality, and 2) solution optimality. Detailed results are in Appendix E. The conclusion is that though the task is simple, $\mathcal{MP}$ helps improve results for monitoring execution, validating results, and suggesting algorithmic approaches.
Thanksgiving Dinner Planning
This problem, detailed in Section 3, evaluates $\mathcal{MP}$’s ability to generate workflows $\mathbf{W^{*}}$ with enhanced constraints and dependencies. It involves multiple agents performing interdependent tasks, representing a complex multi-agent scenario.
Planning performance is compared across three configurations: DeepSeek + $\mathcal{MP}$, GPT4o + $\mathcal{MP}$, and Claude + $\mathcal{MP}$. Metrics include:
@@ 5 Evaluating $\mathcal{MP}$’s Planning Capabilities
5.1 Meta-Planning for Thanksgiving Event
@@
The planner monitors nodes and edges, enabling dynamic adjustments. Full specifications are in Appendix A.
@@
In reactive planning, $\mathcal{MP}$ integrates an early information agent to detect flight delays at departure, enabling timely workflow updates and demonstrating adaptability.
5.2 Experimental Results
@@
As illustrated in Table 6, DeepSeek demonstrated superior scheduling efficiency by optimizing James’s airport wait time for Emily’s pickup, requiring only two iterations. While GPT4o eventually produced a valid solution in three iterations, it created suboptimal travel patterns by having Michael make separate trips. Claude’s solution, though feasible in two iterations, included unnecessary travel between pickup tasks. This experiment highlighted how $\mathcal{MP}$’s explicit constraint specification and common-sense augmentation enabled consistent performance improvement across different LLMs.
@@
GPT4o failed to produce a valid solution, violating three critical constraints without the ability to recognize these violations.
Reactive Planning Performance
The flight delay scenario revealed significant differences between LLMs’ capabilities. DeepSeek demonstrated superior spatial reasoning by routing Michael directly to the airport, an insight that should have come from $\mathcal{MP}$’s common-sense spatial reasoning. This unexpected ability to improve workflow highlights the synergy between $\mathcal{MP}$ and LLM —$\mathcal{MP}$ provided early alert through its information agent (Table 15 in Appendix A.9), which DeepSeek leveraged for immediate replanning.
In contrast, Claude produced two feasible plans but missed the 10:00 AM alert in $\mathbf{W^{R}}$, starting its schedule at 1:00 PM and missing opportunities for proactive actions like early Grandma pickup to free Sarah as a driver. GPT4o failed entirely, producing three constraint violations it couldn’t recognize, preventing further improvements.
Table 7: Reactive Planning Performance (Flight Delay)
@@
GPT4o Failed Failed to maintain critical constraints; unable to recover
@@
6 Conclusions with Impact Statements
Current artificial intelligence systems excel at pattern matching but struggle with deliberate analytical reasoning required for complex planning. While Large Language Models (LLMs) demonstrate impressive capabilities, they face three critical limitations: inability to validate their own outputs (paralleling Gödel’s incompleteness theorems), inconsistent maintenance of constraints during planning, and inadequate integration of common-sense knowledge. Our experiments demonstrate that existing multi-agent systems, focused primarily on coordination, fail to address these limitations.
MACI (Multi-Agent Collaborative Intelligence) advances the state of the art through three key contributions:
@@ 6 Conclusions with Impact Statements
    Meta-Planning Architecture: MACI generates specialized planning workflows by:
    • Constructing role-node and dependency networks.
    • Identifying implicit constraints through common-sense analysis.
    • Establishing validation across distributed agents.
@@
    Workflow Enhancement: As demonstrated in our experiments:
    • Sequential Planning: Enhanced workflow $\mathbf{W^{*}}$ enabled all LLMs to generate feasible solutions.
    • Reactive Planning: Early information agent in $\mathbf{W^{R}}$ enabled proactive replanning.
    • Unexpected synergy: DeepSeek contributed spatial optimizations beyond MACI’s specifications.
@@
    Empirical Validation: Our Thanksgiving dinner planning case study shows:
    • Improving constraint satisfaction on LLMs.
    • Effective handling of crises through early detection.
    • Clear performance differences between LLMs in utilizing MACI’s capabilities.
MACI demonstrates that addressing LLM limitations requires more than simple agent coordination. It demands a structured meta-planning architecture that enables comprehensive constraint awareness and validation. The framework’s success in complex temporal planning, coupled with unexpected improvements from LLM contributions, suggests a promising direction for developing more capable AI planning systems.
Appendices
Appendix A Experiment Details: Meta-Planning for the Thanksgiving Dinner Task

### v2

@@ 4.2.4 Agent Roles in State Space Management
Common agents are designed to handle foundational planning tasks that align with MACI’s state space dimensions (Who, Where, When, What, Why). These agents provide general-purpose functionality that ensures consistency, feasibility, and robustness across planning tasks. Their primary responsibilities include:
  • Constraint Validation Agents: Ensure adherence to temporal, spatial, and resource constraints by verifying the feasibility of generated plans.
  • Common Sense Integration Agents: Identify implicit constraints that may be overlooked, such as transition times, dependencies, or practical limitations.
  • Adaptation Agents: Dynamically adjust plans in response to changes in task environments, such as resource delays or evolving requirements.
  • Performance Evaluation Agents: Assess the quality and efficiency of proposed plans relative to predefined metrics, ensuring continuous improvement.
By addressing these tasks, common agents form the backbone of MACI’s planning architecture. Their modular design enables reuse across multiple domains, and their collaborative functionality ensures they work seamlessly with specialized agents to maintain global consistency and coherence within the planning workflow.
Task-Specific Agents
These agents cater to domain-specific requirements, including task-dependent data and knowledge augmentation, selecting and optimizing planning algorithms, safety and ethics assessment, and emergency response optimization. By leveraging domain expertise, specialized agents extend the capabilities of common agents, enabling MACI to address specialized planning challenges with precision and adaptability.
4.2.5 Seamless Integration and Scalability
The repository’s standardized agent specifications and matching mechanism enable MACI to scale efficiently across domains. By leveraging modular designs and protocol buffers, new agents can be integrated seamlessly into existing workflows, ensuring adaptability and extensibility without compromising performance or consistency.
@@ 4.3 Meta-Planner: Planning a Planner to Plan
The mission of the meta-planner $\mathcal{MP}$ is to construct a planner that generates an actionable workflow for a given task. It does so by analyzing task objectives, identifying roles and constraints, and organizing agents into a structured execution plan. This three-phase approach ensures that every agent and dependency is optimally placed, refined, and validated, leading to robust, task-specific workflows.
@@ 4.3.1 The Meta-Planner Algorithm
The meta-planner operates as a higher-order planning system that formulates workflows as directed graphs:
$\mathbf{W}=(\mathcal{N},\mathcal{E}),\text{ where }\mathcal{N}=A^{*}_{n},\mathcal{E}=A^{*}_{e}.$ (1)
Here, $\mathcal{N}$ denotes roles assigned to agents, and $\mathcal{E}$ represents dependencies between roles, including constraints such as timing, data flow, and supervision requirements.
@@
  1. Extract roles $\mathcal{N}$ from $\mathcal{O}$ (Eq. 2)
  2. Identify dependencies $\mathcal{E}$ from $\mathcal{C_{E}}$ (Eq. 3)
@@
  3. Assign agents to nodes: $\forall n\in\mathcal{N}$, select $\alpha_{n}\in\mathbf{A_{n}}$ (Eq. 4)
  4. Assign agents to dependencies: $\forall e_{ij}\in\mathcal{E}$, select $\alpha_{ij}\in\mathbf{A_{e}}$ (Eq. 5)
@@
        Update role-person mappings $f_{\text{role}}(n,\mathcal{P})$
@@
  return $\mathbf{W^{*}}=\mathbf{W_{current}}$
4.3.2 Meta-Planning Design Elements
@@
The meta-planner extracts roles from task objectives and maps them to required qualifications:
@@
where $n_{i}$ represents a role and $q_{i}$ its required qualifications (e.g., a driver requires a license, a cook requires experience).
@@
Constraints govern role interactions and dependencies. The framework maintains a global constraint set:
@@
where $C_{E}$ represents explicit constraints from task specifications, $C_{I}$ denotes implicit constraints identified by common sense agents, and $C_{D}$ represents derived constraints from agent interactions.
@@
Two categories of agents are assigned based on task requirements:
  • Node Agents (Role Execution):
@@
    These agents are responsible for fulfilling role qualifications and managing people-role assignments.
  • Edge Agents (Dependency Management):
@@
    These agents ensure dependencies between roles are correctly maintained, such as time constraints, spatial relations, and supervisory requirements.
@@ 4.4 Workflow Execution Framework
The final workflow $\mathbf{W^{*}}$ must be executed in a runtime environment. In this work, we evaluate $\mathbf{W^{*}}$ by entering it into an LLM (e.g., GPT4o) alongside the problem statement. A key limitation is that the feedback loop for refining $\mathbf{W^{*}}$ is currently manual, requiring iterative adjustments to optimize execution. Future research will focus on automating this process to enhance adaptability and efficiency.
5 Evaluating $\mathcal{MP}$ vs. Independent LLMs
To assess $\mathcal{MP}$’s performance and adaptability, we adopted a dual-approach experimental structure. The first experiment uses the Traveling Salesperson Problem (TSP) to validate $\mathcal{MP}$’s optimization capabilities. The second involves the Thanksgiving Dinner Planning problem, showcasing $\mathcal{MP}$’s ability to handle complex, real-world challenges with cross-thread dependencies and dynamic adaptability. Due to space constraints, detailed results for these experiments are provided in Appendices D and E, respectively.
5.1 Traveling Salesperson Problem (TSP)
The TSP experiment benchmarks $\mathcal{MP}$ against standalone planners (Claude, DeepSeek, GPT-4o) and their $\mathcal{MP}$-integrated counterparts. The metrics include solution quality and optimality.
Result Summary
Without $\mathcal{MP}$, DeepSeek performs best, while Claude and GPT-4o struggle, each exceeding the optimal travel time by more than $10\%$. With $\mathcal{MP}$, Claude requires two iterations to reach the optimal distance, while both GPT-4o and DeepSeek solve the problem in a single attempt.
Although TSP involves a straightforward, single-thread planning process, $\mathcal{MP}$ still provides notable enhancements. Again, see Appendix D for details.
5.2 Thanksgiving Dinner Planning
This task, detailed in Section 3, evaluates $\mathcal{MP}$’s ability to generate workflows $\mathbf{W^{*}}$ with enhanced constraint and dependency management in the MACI setting. Unlike TSP, this problem involves multiple interdependent agents, introducing complex coordination challenges.
Planning performance is assessed across three configurations: DeepSeek + $\mathcal{MP}$, GPT-4o + $\mathcal{MP}$, and Claude + $\mathcal{MP}$. The prior results in Section 3 show that all LLMs fail the task when executed independently.
Evaluation metrics include:
@@ 5.2 Thanksgiving Dinner Planning
5.2.1 Meta-Planning for Thanksgiving Event
@@
The planner monitors nodes and edges, enabling dynamic adjustments. The full specifications are in Appendix E.
@@
In reactive planning, $\mathcal{MP}$ integrates an alert agent to detect flight delays at departure, enabling timely workflow updates and demonstrating adaptability.
5.2.2 Experimental Results
@@
Table 4 summarizes the detailed schedules documented in Tables 12, 13, and 14, in Appendix E.8. DeepSeek demonstrated superior scheduling efficiency by optimizing James’s airport wait time for Emily’s pickup, requiring only two iterations. While GPT4o eventually produced a valid solution in three iterations, it created suboptimal travel patterns by having Michael make separate trips. Claude’s solution, though feasible in two iterations, included unnecessary travel between pickup tasks. This experiment highlighted how $\mathcal{MP}$’s explicit constraint specification and common-sense augmentation enabled consistent performance improvement across different LLMs.
@@
The flight delay scenario revealed significant differences between LLMs’ capabilities. DeepSeek demonstrated superior spatial reasoning by routing Michael directly to the airport, an insight that should have come from $\mathcal{MP}$’s common-sense spatial reasoning. This unexpected ability to improve workflow highlights the synergy between $\mathcal{MP}$ and LLM —$\mathcal{MP}$ provided early alert through its information agent (Table 15 in Appendix E.9).
Table 5 summarizes the detailed schedules documented in Tables 16, 18, and 19, in Appendix E.9. DeepSeek leveraged the early alert at 10:00 AM for immediate replanning. In contrast, Claude produced two feasible plans but missed the 10:00 AM alert in $\mathbf{W^{R}}$, starting its schedule at 1:00 PM and missing opportunities for proactive actions like early Grandma pickup to free Sarah as a driver. GPT4o failed entirely, producing three constraint violations it could not recognize, preventing further improvements.
Table 5: Reactive Planning Performance (Alert: flight delay)
@@
GPT4o X Failed to maintain critical constraints; unable to recover
@@
6 Conclusion with Impact Statements
This research introduces Multi-Agent Collaborative Intelligence (MACI), a framework designed to overcome fundamental limitations in current LLMs for reasoning and planning. By reimagining computational intelligence through a hierarchical, distributed architecture, MACI represents a structured approach to complex problem-solving.
To validate MACI’s capabilities, two experiments tested its effectiveness in adaptive planning and constraint management. The first experiment, based on the Traveling Salesperson Problem (TSP), demonstrated MACI’s ability to generate globally optimized workflows that guide LLMs in solving this classical problem. The second, a Thanksgiving Dinner Planning task, highlighted MACI’s ability to resolve intricate cross-thread dependencies and dynamically adjust to evolving constraints.
MACI introduces three fundamental innovations that redefine computational reasoning:
Structured Meta-Planning for Constraint-Aware Execution:
The framework implements a structured meta-planner that explicitly separates planning from execution. This architectural approach dismantles the attention-driven limitations that have historically constrained LLM performance, enabling a more deliberate and controlled planning mechanism.
Distributed Validation for Reliable Decision-Making:
MACI establishes a distributed validation mechanism that enhances system reliability. By incorporating independent verification agents, the framework transcends the probabilistic limitations of traditional language models, introducing a new paradigm of self-checking computational intelligence.
Proactive Multi-Agent Coordination for Adaptive Planning:
The system enables a proactive approach to multi-agent coordination. Unlike previous static collaboration frameworks, MACI facilitates real-time constraint resolution and dynamic replanning, significantly advancing adaptive computational reasoning.
By redefining intelligent planning, MACI lays the foundation for AI systems that operate with greater precision, adaptability, and contextual awareness, bridging the gap between static LLM reasoning and dynamic real-world decision-making.
Appendices
Appendix A Validation and Recovery Protocols
$\text{recover}(s_{t},s_{\text{failed}})\rightarrow s_{\text{valid}}$ (7)
A.1 Operations Research Techniques in Validation Protocols
Integration in Multi-Agent Systems
Appendix B $\mathsf{MACI}$ Additional Design Considerations
B.1 Cross-Domain Generalization
B.2 Dynamic Agent Registration and Evolution
B.3 Scalability and Resource Efficiency
B.4 Empirical Evaluation Across Domains
B.5 Challenges and Future Directions
Appendix C Additional Tables and Figures
C.1 Notation
Table 6 presents all symbols used throughout this paper.
Table 6: Symbol Definitions
Symbol Definition Symbol Definition
Basic Sets
$\mathcal{O}$ Planning objectives $\mathcal{P}$ Available people
$\mathcal{M}$ Performance metrics $\mathcal{Q}$ Role qualifications
Workflow Components
$\mathbf{W}$ Workflow network $\mathcal{N}$ Roles (nodes)
$\mathcal{E}$ Dependencies (edges) $\mathbf{A}$ Agent repository
$n_{i}$ Individual role $e_{ij}$ Role dependency
Agent Functions
$f_{\text{role}}$ Role-agent mapping $f_{\text{edge}}$ Edge-agent mapping
Optimization
$A^{*}_{n}$ Selected node agents $A^{*}_{e}$ Selected edge agents
$\mathbf{W^{*}}$ Optimal workflow $V^{*}$ Best validation score
C.2 Example Common Agents
@@
  4. 4.
  5. 5.
  6. 6.
  7. 7.
  8. 8.
  9. 9.
  10. 10.
Appendix D Traveling Salesman Problem Experiment
D.1 General Problem Specification
  • Inputs: N locations, distance matrix D[N][N]
  • Output: Optimal tour T minimizing total distance
  • Constraints: Each location visited once, return to start
Computational Complexity - Brute Force
For N locations:
  • Number of possible tours = (N-1)!/2
  • Time complexity = O(N!)
  • Space complexity = O(N2)
Solution Methods
  1. 1.
  2. 2.
  3. 3.
D.2 $\mathbf{W^{*}}$: MACI Generated Planner for TSP
Node Components (N)
For TSP with n locations:
$N=\{n_{\text{route}},n_{\text{dist}},n_{\text{valid}}\},\text{ where}$ (8)
    $n_{\text{route}}$: Route generation role
    $n_{\text{dist}}$: Distance calculation role
    $n_{\text{valid}}$: Solution validation role
Edge Dependencies (E)
    $e_{\text{spatial}}$: Distance constraints between locations
    $e_{\text{sequence}}$: Visit order constraints
    $e_{\text{complete}}$: Tour completion requirements
Agent Assignments
Node Agents ($A_{n}$):
  • Route Generation Agent: Generates candidate tours
  • Distance Calculator Agent: Computes tour lengths
  • Solution Validator Agent: Verifies tour validity
Edge Agents ($A_{e}$):
  • Spatial Constraint Agent: Monitors distance feasibility
  • Sequence Monitor Agent: Ensures valid visit order
  • Completion Checker Agent: Verifies tour completeness
Algorithm Selection
Validation Function
D.3 Experiments, From Small to Large N
  1. 1.
    N=5: Establish ground truth via brute force
  2. 2.
    N=10,20,100: Test LLM performance degradation
  3. 3.
    Metrics:
    • Solution quality vs optimal
    • Computation attempts before giving up
    • Error recognition capability
D.3.1 Small Campus Tour (N=5)
Plan an optimal route for campus tour guide visiting 5 key locations:
  • A: Admissions Office (start/end)
  • B: Library
  • C: Student Center
  • D: Science Building
  • E: Sports Complex
Distance Matrix (minutes)
Constraints
  • Tour starts/ends at Admissions (A)
  • Each location visited exactly once
  • Total possible routes: (5-1)!/2 = 12
  • Optimal solution can be verified by hand
3.1.1. $\mathbf{W^{*}}$ Workflow Components
Node Components (N)
  • $n_{\text{route}}$: Generates permutations starting/ending at A.
  • $n_{\text{dist}}$: Computes tour length using distance matrix $D$.
  • $n_{\text{valid}}$: Checks tour validity (start/end at A, no repeats).
Edge Dependencies (E)
  • $e_{\text{spatial}}$: Enforces distance constraints from $D$.
  • $e_{\text{sequence}}$: Ensures visit order consistency.
  • $e_{\text{complete}}$: Validates all 5 locations are visited.
Agent Assignments
  • Node Agents:
    • Route Generation Agent (handles $n_{\text{route}}$)
    • Distance Calculator Agent (handles $n_{\text{dist}}$)
    • Solution Validator Agent (handles $n_{\text{valid}}$)
  • Edge Agents:
    • Spatial Constraint Agent (enforces $e_{\text{spatial}}$)
    • Sequence Monitor Agent (enforces $e_{\text{sequence}}$)
    • Completion Checker Agent (enforces $e_{\text{complete}}$)
Selected Algorithm
Brute-force.
Validation Function
3.1.2. Solution Steps
Step 1: Problem Parsing
  • Input: 5 locations with distance matrix $D$.
  • Initialize node/edge agents and constraints.
Step 2: Route Generation ($n_{\text{route}}$ Agent)
  • Generate all valid permutations: $\frac{(5-1)!}{2}=12$ routes.
  • Example permutations:
Step 3: Distance Calculation ($n_{\text{dist}}$ Agent)
  • Compute total time for each route using $D$.
Step 4: Solution Validation ($n_{\text{valid}}$ Agent)
  • Check all routes for:
    • No duplicate visits.
Step 5: Edge Agent Validation
  • Spatial Constraint Agent: Verify $D_{i,j}$ matches edge weights.
Step 6: Apply Validation Function
  • Assign $V=-\infty$ to invalid routes.
[…]
  • Identify minimal $V=-24$ (i.e., maximal tour length 24 mins).
3.1.3 Solution
Optimal tour time: 242424 minutes, achieved by three routes:
D.3.2 Large Campus Tour (N=10)
Plan an optimal route for a guided tour through 10 locations:
  • Distance Matrix: Asymmetric travel times (minutes)
Planner Best Results Algorithm Iters. Advantages Limitations
Algorithm Selection
3.2.1. ACO Method
Parameters
  • 100 ants, 50 iterations, and $\rho=0.1$ evaporation
  • $\alpha=1$ (pheromone weight), and $\beta=2$ (heuristic weight)
Algorithm
1:  Initialize $\tau_{ij}\leftarrow 1.0$, $\eta_{ij}\leftarrow 1/D_{ij}$
2:  for 50 iterations do
3:     for all 100 ants do
5:        Record tour length $L_{k}$
6:     end for
7:     Evaporate pheromones: $\tau_{ij}\leftarrow 0.9\tau_{ij}$
8:     Deposit pheromones: $\tau_{ij}\leftarrow\tau_{ij}+\sum\frac{10}{L_{k}}$
9:     Track best tour
10:  end for
3.2.2. Performance Comparison
D.4 TSP Experiment Conclusion
Appendix E Experiment Details: Meta-Planning for the Thanksgiving Dinner Task